OpenStudy (anonymous):

How to find a1 if r=-4, an=3072 and Sum of ak=2457 k=1 n

OpenStudy (amistre64):

do you know the formula for the sum of a geometric sequence?

OpenStudy (anonymous):

Yes, I more need help finding n

OpenStudy (anonymous):

\[\sum_{k=1}^{n} a _{k}=2457\]

OpenStudy (anonymous):

how do I find n?

OpenStudy (amistre64):

the formula, if you may have forgotten is:\[\sum_{1}^{n}a_k=a_1\frac{1-r^n}{1-r}\] yeah, finding n may be tricky

OpenStudy (amistre64):

we might have to define n in terms of a1 ... maybe

OpenStudy (anonymous):

a1 is the goal to find

OpenStudy (amistre64):

a1 + a1*r + a1*r^2 + .... + a1 r^(n-1) ; since we started with r^0 a1 r^(n-1) = an

OpenStudy (amistre64):

we have 2 equations and 2 unknowns

OpenStudy (anonymous):

r and \[a_{n}\] are kown

OpenStudy (amistre64):

\[a_1r^{n-1}=3072\] \[a_1\frac{1-r^n}{1-r}=2457\] got a new keyboard and the put the \ key in the wrong spot .... or at least a different spot and its really annoying me lol